OpenStudy (mathmath333):

Brass is an alloy of copper and zinc. Bronze is an alloy containing 80% copper, 4% of zinc and 16% of tin. A fused mass of brass and bronze is found to contain 74% of copper, 16% zinc, and 10% of tin. The ratio of copper to zinc in brass is? a.) 64% and 36% b.) 33% and 67% c.) 50% and 75% d.)35% and 65%

OpenStudy (mathmath333):

@amistre64

OpenStudy (amistre64):

if this is possible to solve, id say we simply remove the bronze ratio to determine whats left

OpenStudy (mathmath333):

can u elaborate

OpenStudy (amistre64):

we are given that bronze has a given ratio pf parts, and that the combined mass contains bronze in it, remove the bronze part by an appropriate ratio

OpenStudy (amistre64):

since we have no tin in brass, what is the proper scalar to get .16t = .10 ?

OpenStudy (mathmath333):

t=5/8

OpenStudy (amistre64):

then 5/8 of the copper is bronze, and 5/8 of the zinc is bronze remove them to find whats left is what comes to mind

OpenStudy (mathmath333):

u mean 74-5/8(74) like this ?

OpenStudy (amistre64):

.74 * 5/8 = bronze so .74 * 3/8 = brass copper and likewise, .16*3/8 should give us brass zinc

OpenStudy (amistre64):

74 - 5/8(74) seems fair as well. took some time to process what you meant :)

OpenStudy (mathmath333):

0.74*3/8=0.277

OpenStudy (mathmath333):

so how do i convert 0.277 in %

OpenStudy (amistre64):

.277 is just how much copper we have left over, how much zinc is left over?

OpenStudy (mathmath333):

zince 0.16*3/8=0.06

OpenStudy (amistre64):

Bronze contains 80% copper 4% of zinc 16% of tin A fused mass of brass and bronze is found to contain .74 of copper = .80 bronze + x brass .16 of zinc = .04 bronze + y brass .10 of tin = .16 bronze + 0 brass the tin row can solve for bronze, 5/8 .74 of copper = .80*5/8 + x brass .16 of zinc = .04*5/8 + y brass .10 of tin = .16*5/8 + 0 brass now we have 2 equations in 2 unknowns

OpenStudy (amistre64):

something along these lines should work out

OpenStudy (mathmath333):

book gives answer 64% copper and 36% zinc

OpenStudy (amistre64):

thats what i get as well

OpenStudy (amistre64):

.74 = .80*5/8 + x*3/8 .16 = .04*5/8 + y*3/8 .10 = .16*5/8 x and y are both solvable

OpenStudy (mathmath333):

thanks very much